OpenStudy (anonymous):

Rod brought 40 juice bottles to the picnic. He brought apple juice bottles and the rest were orange juice bottles. What was the ratio of apple juice bottles to orange juice bottles? A. 2 to 1 B. 3 to 10 C. 7 to 3 D. 7 to 10

OpenStudy (anonymous):

Can anyone help me solve will fan and medal

OpenStudy (anonymous):

how many bottles of applejuice?

OpenStudy (anonymous):

28

OpenStudy (anonymous):

so there are 28 bottles of A.J. and 12 bottles of O.J.

OpenStudy (anonymous):

@SkaterBoyShawn

OpenStudy (anonymous):

@Deadpool117

OpenStudy (akashdeepdeb):

Total bottles = 40 Apple juice bottles = 28 Orange bottles = 40 - 28 = 12 Ratio of Apple Juice Bottles to Orange bottles = \(\frac{Apple}{Orange} =~ ?\)

OpenStudy (anonymous):

@gabbyalicorn @HannahA @Deadpool117 @Deadpool25

OpenStudy (gabbyalicorn):

Sorry im not the best at math.

OpenStudy (anonymous):

@jane11509 @mathmate @mathscience4ever

OpenStudy (mathmate):

@Ashleyistheboss2 It's not the best OS etiquette to keep tagging people when people are leading you to the answer. I suggest you work with the first person who helps you to get your answer fast.